The price of coffee increase from rs 50 per kg to rs 70 per kg and as a result the demands for tea increased from 5 kg to 10 kg what is cross elasticity of demand of tea for coffee? What is the relationship between tea and coffee


1
Expert's answer
2021-12-20T10:20:27-0500

Cross Elasticity of demand for tea = "\frac{\%\;change\; in\; quantity\; demanded\; for\; tea\; }{\%\; change\; in\; price\; of\; coffee\; }"


"=\frac{Q_{2} -Q_{1}}{(Q_{2}+Q_{1})/2 } \div \frac{P_{2} -P_{1}}{(P_{2}+P_{1})/2 }"


"=\frac{10 -5}{(10+5)/2 } \div \frac{70 -50}{(70+50)/2 }"


"=\frac{2}{3}\div\frac{1}{3}=2"

Tea is a substitute to coffee since the cross price elasticity of demand is positive